The vector having a magnitude of 10 and perpendicular to the vevtor 3i^-4j^ is??

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Give the vector we a chance to need to discover is A=ai+bj 
let the vector given is B=3i-4j 
for 2 vectors to be opposite the spot item should be 0. 
A.B=0 
=>(ai+bj).(3i-4j)=0 
=>3a-4b=0 
=>b=3a/4 
|A|=√(a^2+b^2 ) 
=>a2+b2=100 
=>a2+9a2/16=100 
=>a=+-8 
b=+-6 
A=8i+6j or - 8i-6j or 8i-6j or - 8i+6j 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)